# Deepfakes, Phrenology, Surveillance, and More! A Taxonomy of AI Privacy Risks

### v1

@@ 4.2.1. Identification (124/321)
Scale (20/124): AI technologies enabled automated identification at scale. One example is Facebook’s now-disabled Tag Suggestions product, through which Facebook demonstrated its ability to automatically identify individuals from uploaded photos. When this feature was in use, Facebook had 1.4 billion daily active users555https://investor.fb.com/home/default.aspx; still, “any time someone uploads a photo that includes what Facebook thinks is your face, you’ll be notified even if you weren’t tagged” (Simonite, 2017).
Latency (24/124): Once the models are trained, AI technologies also allow identification risks to occur more quickly, in nigh real-time. For example, in 2019, the Italian government was on the verge of implementing a real-time facial recognition system across football stadiums that “prevent individuals who are banned from sports competitions from entering stadiums.” It also picked up audiences’ “racist conversations” to alert law enforcement authorities to the presence of racist fans (Spielkamp, 2019).
Robustness (7/124): AI also allows for robust identification even with low-quality data. Clearview AI, a facial recognition application that aids U.S. law enforcement in identifying wanted individuals, claims to be able to identify people under a range of obfuscation conditions: “[a] person can be wearing a hat or glasses, or it can be a profile shot or partial view of their face” (Hill, 2020). Similarly, models trained on Simulated Masked Face Recognition Dataset (SMFRD)666https://github.com/X-zhangyang/Real-World-Masked-Face-Dataset#download-datasets are capable of identifying persons with a mask on, “violating the privacy of those who wish to conceal their face” (Wiggers, 2021).
Ubiquity (73/124): Finally, AI technologies can enable identification risks in situated physical environments such as public places (e.g., (Mozur, 2018)), stores (e.g., (Hodgson, 2019)), and classrooms (e.g.,(Reuters, 2018)). For example, XPeng Motors, a Chinese electric vehicle firm, was reported for using facial recognition-embedded cameras to collect biometric data of customers visiting their stores (Global Times, 2021).
@@ 4.2.2. Aggregation (49/321)
Scale (23/49): One of the unique strengths of AI systems is that they automate complex processes into simple programs that overcome human limitations. While controversial, many public sectors still utilize algorithmic tools in high-stake contexts such as social work (Kawakami et al., 2022) and services for the unhoused (Kuo et al., 2023) to prioritize limited resources. AI technologies can process vast amounts of personal data to infer invasive things about individuals not explicit in those data. For example, an AI start-up created a service that assesses a prospective babysitter’s likelihood to engage in risky behaviors such as drug abuse and bullying by “scan[ning] … thousands of Facebook, Twitter and Instagram posts” (Harwell, 2018). Similarly, a recruiting technology firm created a system to assess the abilities of prospective employees based on an assemblage of data sources, including written answers, behavior, intonation, and speech (Harwell, 2019).
Latency (11/49): AI technologies can learn how to perform complicated detection and identification tasks nigh instantly. Technologies have been developed to estimate employee performance in-the-moment (Wakefield, 2021a), and to forecast what one might write in emails (Tsukayama, 2021). AI technologies have also been developed to predict when end-users might be ovulating (Hunter and Kelly, 2022), and their moment-to-moment risk of committing suicide (Beckett, 2019).
Ubiquity (5/49): AI technologies can make physical objects and environments smarter and more responsive. Smart home devices, for example, allow for automated control of home appliances, dynamic temperature control to strike an optimal balance between energy consumption and comfort, and voice user interfaces (Alam et al., 2012). These features require AI technologies to continuously monitor data streamed from physical sensors, creating new types of aggregation risks in situated environments. For example, smart speaker microphone feeds have been used to infer who is present in a room, who is speaking, and other information that can be algorithmically inferred from voice data (noa, 2018).
Prediction (10/49): Finally, AI technologies are able to robustly forecast future behaviors and states. This forecasting can be used, for example, to help proactively identify health risks, plan optimal routes to avoid predictable traffic, and estimate retirement savings. These capabilities also create a new type of predictive aggregation risk. For example, in 2018, Argentina’s government deployed an AI model that predicted teen pregnancy in low-income areas from their first name, last name, and address (Jemio et al., 2022). AI has also been used for crime prediction. For example, in 2018, law enforcement in the United Kingdom aimed to predict serious violent crime using AI: “[t]he system, called the National Data Analytics Solution, uses a combination of AI and statistics to try to assess the risk of someone committing or becoming a victim of gun or knife crime” (Baraniuk, 2018).
@@ 4.2.3. Phrenology / Physiognomy (27/321)
Phrenology and Physiognomy are debunked pseudosciences that postulated that it is possible to make reliable inferences about a person’s personality, character, or predispositions from an analysis of their outer appearance and/or physical characteristics (noa, 2012). Prior work warned that irresponsible use of AI classification models to make spurious inferences about people from inputs that capture their outer appearance could usher in a revival of these pseudosciences (Arcas, 2017). Our analysis reveals that AI technologies are indeed being used in this way, resulting in a new category of privacy risk not captured by Solove’s initial taxonomy. We define phrenology/physiognomy risks as the use of AI to infer personality, social, and emotional attributes about an individual from their physical attributes. This risk stems from AI’s ability to learn correlations between arbitrary inputs (e.g., images, voices) and outputs (e.g., one’s demographic information).
@@ 4.2.5. Exclusion (149/321)
Prior work has shown that it can be challenging to ensure agency to any individual over their data regarding how data they have shared online can and cannot be used by such models (Obar, 2020), and that it can be deliberately made complex for individuals to remove their data from the dataset (Birhane et al., 2021). Additionally, when commercial AI models are “black boxes,” the general public has no means to audit how personal data is used by AI (e.g., Clearview AI).
@@ 4.3.1. Exposure (17/321)
Exposure risks encompass revealing sensitive private information that people view as deeply primordial that we have been socialized into concealing (Solove, 2006). Traditionally, these risks arise when an individual’s private activities are recorded and disseminated to others without consent. AI technologies can create new types of exposure risks via generative techniques that can create, reconstruct, manipulate content (i.e., deepfake techniques) (10/17) and expose inferred sensitive end-user attributes predicted by AI/ML (7/17) (e.g., one’s interests (Levin, 2017b)).
@@ 4.3.2. Distortion (20/321)
Some models can generate realistic audio of individuals. For example, Prime Voice AI, a text-to-voice generator, was misused to create the voices of celebrities to “make racist remarks about Alexandria Ocasio-Cortez (the US House representative)”, and that the AI-generated clips “run the gamut from harmless, to violent, to transphobic, to homophobic, to racist.” (Cox, 2023; Harrison, [n. d.]). Other AI-created distortion risks are less egregious, but raise important questions about expectations around privacy in light of how generative AI can be used to simulate the likeness of those who have passed. For example, the filmmaker of a documentary was revealed to be using deepfake technology to create scenes, with the likeness of an actor who had passed away, for lines “he wanted [Anthony] Bourdain’s (the main character of the documentary) voice for but had no recordings of” (Leon, 2021).
@@ 4.4.1. Intrusion (160/321)
Intrusion risks encompass actions that disturb one’s solitude in physical space (Solove, 2006). For six of the 160 intrusion incidents we identified, we noted that the AI technologies described in the incident did not fundamentally change the risk described in the incident: the intrusion would have remained as described even without the capabilities and/or requirements of AI. One example is the use of digital screens in stores to show customers personalized ads (Meyersohn, 2022): the intrusion would remain, even if the system did not use AI. However, we identified two ways AI can exacerbate intrusion risks, i.e., by increasing scale and ubiquity.
Scale (113/160): The capabilities of AI technologies (e.g., to identify a person and detect behaviors) enable a centralized surveillance infrastructure that creates large-scale intrusion risks; the requirements of AI (e.g., access to vast troves of data and GPU servers) necessitate this infrastructure. For example, Pharmaceutical University in Nanjing, China, implemented a recognition system at various locations on campus to closely monitor students’ attendance and learning behaviors (Tone, 2019; Zhang, 2019). Similarly, employers are increasingly incorporating AI-infused workplace monitoring technologies that collect data from employees’ smartwatches (The Local, 2020) and computer webcams (Walker, 2021) to track their performance, absence, and time-on-task.
Ubiquity (41/160): The capabilities of AI can also turn everyday products such as doorbells and wristbands into powerful nodes in a ubiquitous surveillance infrastructure. For example, Ring, a smart doorbell that enables homeowners to monitor activities and conversations near where the doorbell is installed, has raised concern due to “the device’s excessive ability” to capture data of an individual’s neighbors (Milmo, 2021). Similarly, Amazon’s Halo fitness tracker uses AI to analyze a user’s conversations to highlight when and how often that user spoke in a manner that was indicative of their being “happy, discouraged, or skeptical” (O’Hara, 2021).
@@ 5. Discussion
Our findings demonstrate the many ways modern advances in AI meaningfully change privacy risks relative to Solove’s widely cited taxonomy of privacy (Solove, 2006). Across the 321 AI privacy incidents we analyzed, roughly 7% of the cases did not involve privacy risks that were created or exacerbated by AI. For example, we encountered instances where a product that happened to include AI was subject to a data breach in which users’ personal data was compromised (Abrams, 2020). Nevertheless, in approximately 93% of the cases we analyzed, the unique capabilities and data requirements of the AI technologies involved in the incident either created a new type of privacy risk, or exacerbated a known risk.
@@
We found that the unique capabilities of AI create new types of privacy risks. For example, AI creates new data processing risks in its ability to identify the activity of individuals even with low-quality data, and in its ability to forecast future outcomes. AI creates a new category of phrenology/physiognomy risks by enabling the creation of spurious classifiers correlating physical attributes with social, emotional, and personality traits. AI creates new types of data dissemination risks in its ability to generate human-like media, e.g., by generating a realistic fake video of an individual. We also found that the data requirements of AI exacerbate privacy risks we have grappled with for decades. For example, AI technologies can lead to more severe surveillance regarding scale and ubiquity; exacerbate secondary use, exclusion, insecurity, disclosure, and increased accessibility risks in the processing and disseminating of personal data; and intrude upon people’s personal space at a larger scale and with greater pervasiveness than ever before.
@@
Equipped with the knowledge of how AI has changed privacy risks, we first discuss risks documented by Solove’s original taxonomy (Solove, 2006) for which we did not find matching incidents in our incident database. In theory, future advances in and/or the use of AI may entail risks in these categories, so it is worth discussing them as privacy risks that AI may change in the future. We then next discuss a number of ways we expect this taxonomy might be useful for both future research and practice.

### v2

@@ 4.2.1. Identification (124/321)
AI technologies enabled automated identification at scale (20/124). One example is Facebook’s now-disabled Tag Suggestions product, through which Facebook demonstrated its ability to automatically identify individuals from uploaded photos. When this feature was in use, Facebook had 1.4 billion daily active users555https://investor.fb.com/home/default.aspx; still, “any time someone uploads a photo that includes what Facebook thinks is your face, you’ll be notified even if you weren’t tagged” (Simonite, 2017).
AI technologies allow identification risks to occur more quickly, in nigh real-time (24/124), once the models are trained. For example, in 2019, the Italian government was on the verge of implementing a real-time facial recognition system across football stadiums that “prevent individuals who are banned from sports competitions from entering stadiums.” It also picked up audiences’ “racist conversations” to alert law enforcement authorities to the presence of racist fans (Spielkamp, 2019).
In addition, AI technologies allow for robust identification even with low-quality data (7/124). Clearview AI, a facial recognition application that aids U.S. law enforcement in identifying wanted individuals, claims to be able to identify people under a range of obfuscation conditions: “[a] person can be wearing a hat or glasses, or it can be a profile shot or partial view of their face” (Hill, 2020). Similarly, models trained on Simulated Masked Face Recognition Dataset (SMFRD)666https://github.com/X-zhangyang/Real-World-Masked-Face-Dataset#download-datasets are capable of identifying persons with a mask on, “violating the privacy of those who wish to conceal their face” (Wiggers, 2021).
Finally, AI technologies enable ubiquity identification risks in situated physical environments (73/124) like public places (e.g., (Mozur, 2018)), stores (e.g., (Hodgson, 2019)), and classrooms (e.g.,(Reuters, 2018)). For example, XPeng Motors, a Chinese electric vehicle firm, was reported for using facial recognition-embedded cameras in their stores to collect biometric data of customers (Global Times, 2021).
@@ 4.2.2. Aggregation (49/321)
One of the unique strengths of AI systems is that they automate complex processes into simple programs that overcome human limitations. While controversial, many public sectors still utilize algorithmic tools in high-stake contexts such as social work (Kawakami et al., 2022) and services for the unhoused (Kuo et al., 2023) to prioritize limited resources. To that end, AI technologies create aggregation at scale (23/49) by processing vast amounts of personal data to infer invasive things about individuals not explicit in those data. For example, an AI start-up created a service that assesses a prospective babysitter’s likelihood to engage in risky behaviors such as drug abuse and bullying by “scan[ning] … thousands of Facebook, Twitter and Instagram posts” (Harwell, 2018).
AI technologies perform complicated inferencing tasks nigh instantly (11/49). Technologies have been developed to estimate employee performance in-the-moment (Wakefield, 2021a), and to forecast what one might write in emails (Tsukayama, 2021). AI technologies have also been developed to predict when end-users might be ovulating (Hunter and Kelly, 2022), and their moment-to-moment risk of committing suicide (Beckett, 2019).
AI technologies can make physical objects and environments smarter and more responsive, enabling ubiquitous aggregation risks (5/49). Smart home devices, for example, allow for automated control of home appliances, dynamic temperature control to strike an optimal balance between energy consumption and comfort, and voice user interfaces (Alam et al., 2012). These features require AI technologies to continuously monitor data streamed from physical sensors, creating new aggregation risks in situated environments. For example, smart speaker microphone feeds have been used to infer who is present in a room, who is speaking, and other information that can be algorithmically inferred from voice data (noa, 2018).
Finally, AI technologies enable forecasting future behaviors and states based on historical data. This forecasting can be used, for example, to help proactively identify health risks, plan optimal routes to avoid predictable traffic, and estimate retirement savings. These capabilities of AI, however, also create a new type of predictive aggregation risk (10/49). For example, in 2018, Argentina’s government deployed an AI model that predicted teen pregnancy in low-income areas from their first name, last name, and address (Jemio et al., 2022). AI has also been used for crime prediction. For example, in 2018, law enforcement in the United Kingdom aimed to predict serious violent crime using AI based on “records of people being stopped and searched and logs of crimes committed” (Baraniuk, 2018).
@@ 4.2.3. Phrenology / Physiognomy (27/321)
Phrenology and Physiognomy are debunked pseudosciences that postulate that it is possible to make reliable inferences about a person’s personality, character, or predispositions from an analysis of their outer appearance and/or physical characteristics (noa, 2012). Beyond the baseless prediction made from historical data streams discussed in Aggregation risks (Section 4.2.2), phrenology/physiognomy risks pose unique downstream privacy harms distinct from aggregation risks: whereas aggregation risks primarily arise from the collection and combination of disparate pieces of information to make deductive inferences about individuals, phrenology/physiognomy risks introduce new and unfounded inferences about an individual’s internal characteristics (e.g., their preferences and proclivities). Moreover, while aggregation risks generally come from the combination of factual and observable data streams over which users can have some awareness and control (e.g., purchasing habits), phrenology/physiognomy risks arise from making inferences over physical characteristics over which users have no control. Moreover, beyond the harm to the individual, there is also a broader societal harm: prior work has warned that irresponsible use of AI classification models could usher in a revival of these pseudosciences (Arcas, 2017; Stark and Hutson, 2021) by, e.g., motivating surveillance institutions to train AI models to make spurious inferences about a person’s preferences, personality, and character from inputs that capture their outer appearance. Our analysis reveals that AI technologies are indeed being used in this way, resulting in a new category of privacy risk not captured by Solove’s initial taxonomy. We define phrenology/physiognomy risks as the use of AI to infer personality, social, and emotional attributes about an individual from their physical attributes. This risk stems from AI’s ability to learn correlations between arbitrary inputs (e.g., images, voices) and outputs (e.g., one’s demographic information).
@@ 4.2.5. Exclusion (149/321)
Prior work has shown that it can be challenging to ensure agency to any individual over their data regarding how data they have shared online can and cannot be used by such models (Obar, 2020), and that it can be deliberately made complex for individuals to remove their data from the dataset (Birhane et al., 2021). Additionally, when commercial AI models are “black boxes,” the general public has no means to audit how personal data is used by AI (e.g., Clearview AI). Finally, “algorithmic inclusion” — i.e., ensuring that everyone is included in a system — is often seen as a more desirable way to build AI systems in the context of AI ethics. These “inclusive AI” approaches, however, need to be balanced against exclusion-based privacy risks (Albert and Delano, 2022; Andrus and Villeneuve, 2022): when more people’s data are captured to build inclusive systems, those people may be subject to increased exclusion risk if their data is collected without adequate consent and control.
@@ 4.3.1. Exposure (17/321)
Exposure risks encompass revealing sensitive private information that people view as deeply primordial that we have been socialized into concealing (Solove, 2006). Traditionally, these risks arise when an individual’s private activities are recorded and disseminated to others without consent. AI technologies can create new types of exposure risks via generative techniques that can create, reconstruct, manipulate content (i.e., deepfake techniques) (10/17) and expose inferred sensitive end-user attributes predicted by AI/ML (e.g., one’s interests (Levin, 2017b)) (7/17).
@@ 4.3.2. Distortion (20/321)
Some models can generate realistic audio of individuals. For example, Prime Voice AI, a text-to-voice generator, was misused to create the voices of celebrities to “make racist remarks about Alexandria Ocasio-Cortez (the US House representative)”, and that the AI-generated clips “run the gamut from harmless, to violent, to transphobic, to homophobic, to racist.” (Cox, 2023; Harrison, 2023). Other AI-created distortion risks are less egregious, but raise important questions about expectations around privacy in light of how generative AI can be used to simulate the likeness of those who have passed. For example, the filmmaker of a documentary was revealed to be using deepfake technology to create scenes, with the likeness of an actor who had passed away, for lines “he wanted [Anthony] Bourdain’s (the main character of the documentary) voice for but had no recordings of” (Leon, 2021).
@@ 4.4.1. Intrusion (160/321)
Intrusion risks encompass actions that disturb one’s solitude in physical space (Solove, 2006). For six of the 160 intrusion incidents we identified, we noted that the AI technologies described in the incident did not fundamentally change the risk described in the incident: the intrusion would have remained as described even without the capabilities and/or requirements of AI. One example is the use of digital screens in stores to show customers personalized ads (Meyersohn, 2022): the intrusion would remain, even if the system did not use AI. However, we identified two ways AI can exacerbate intrusion risks that increase their scale and ubiquity.
The capabilities of AI technologies (e.g., to identify a person and detect behaviors) enable a centralized surveillance infrastructure that creates large-scale intrusion risks (113/160); the requirements of AI (e.g., access to vast troves of data and GPU servers) necessitate this infrastructure. For example, Pharmaceutical University in Nanjing, China, implemented a recognition system at various locations on campus to closely monitor students’ attendance and learning behaviors (Tone, 2019; Zhang, 2019). Similarly, employers are increasingly incorporating AI-infused workplace monitoring technologies that collect data from employees’ smartwatches (The Local, 2020) and computer webcams (Walker, 2021) to track their performance, absence, and time-on-task.
The capabilities of AI can also turn everyday products (e.g., doorbells, wristbands) into powerful nodes in a ubiquitous surveillance infrastructure (41/160). For example, Ring, a smart doorbell that enables homeowners to monitor activities and conversations near where the doorbell is installed, has raised concern due to “the device’s excessive ability” to capture data of an individual’s neighbors (Milmo, 2021). Similarly, Amazon’s Halo fitness tracker uses AI to analyze a user’s conversations to highlight when and how often that user spoke in a manner that was indicative of their being “happy, discouraged, or skeptical” (O’Hara, 2021).
@@ 5. Discussion
Our findings demonstrate the many ways modern advances in AI meaningfully change privacy risks relative to how we conceived of privacy risks prior to these advances, as captured by Solove’s widely cited taxonomy of privacy (Solove, 2006). Across the 321 AI privacy incidents we analyzed, roughly 7% of the cases did not involve privacy risks that were created or exacerbated by AI. For example, we encountered instances where a product that happened to include AI was subject to a data breach in which users’ personal data was compromised (Abrams, 2020). Nevertheless, in approximately 93% of the cases we analyzed, the unique capabilities and data requirements of the AI technologies involved in the incident either created a new type of privacy risk, or exacerbated a known risk.
@@
We found that the unique capabilities of AI create new types of privacy risks. For example, AI creates new data processing risks in its ability to identify the activity of individuals even with low-quality data, and in its ability to forecast future outcomes. AI creates a new category of phrenology/physiognomy risks by enabling the creation of spurious classifiers correlating physical attributes with social, emotional, and personality traits. AI creates new types of data dissemination risks in its ability to generate human-like media, e.g., by generating a realistic fake video of an individual. We also found that the data requirements of AI exacerbate privacy risks we have grappled with for decades. For example, AI technologies can lead to more pervasive, larger scale surveillance than before; exacerbate secondary use, exclusion, insecurity, disclosure, and increased accessibility risks in the processing and disseminating of personal data; and, increase the ways in which computing can intrude upon people’s personal space.
@@
Equipped with the knowledge of how AI has changed privacy risks, we first discuss how the current AI/ML methods fall short and only address a subset of the AI privacy risks identified in our taxonomy (Section 5.1). Then, we present our taxonomy as a living structure that can be expanded with risks documented by Solove’s original taxonomy (Solove, 2006) in cases where we did not find matching incidents in our incident database (Section 5.2). In theory, future advances in and/or the use of AI may entail risks in these categories, so it is worth discussing them as privacy risks that AI may change in the future. Moreover, we discuss a number of ways we expect this taxonomy might be useful for both future research and practice (Sections 5.1.1 and 5.2.1).
5.1. Charting the design space for privacy-preserving AI/ML work
Our findings broaden the design space for privacy-preserving AI and ML. For example, a recent meta-review of HAI principles and guidelines argues that privacy in ML-driven systems centers around the protection, control, and agency over personal data (Yurrita et al., 2022). Based on our findings, these criteria only consider a small subset of the AI privacy risks we identified: they consider some — but not all — of the data collection and processing risks exacerbated by AI, and do not at all consider the data processing and dissemination risks newly created by AI. In this section, we provide an overview of how the existing tools and approaches, that aim to help practitioners build privacy-preserving AI systems (Yurrita et al., 2022; Meurisch and Mühlhäuser, 2022; Wong et al., 2023), fall short of effectively identifying and addressing many AI privacy risks.
Differential Privacy and Federated Learning
Differential Privacy (DP) (Nasr et al., 2018) and Federated Learning (FL) (Li et al., 2020) are commonly thought of as approaches to “privacy-preserving” machine learning where 1) the model output is insensitive to the presence or absence of data on an individual in a dataset, and 2) the model provider only learns and improves the model in an aggregated manner. Tools such as Diffprivlib131313https://github.com/IBM/differential-privacy-library (Holohan et al., 2019) and IBM Federated Learning141414https://github.com/IBM/federated-learning-lib (Holohan et al., 2019) have been used by practitioners to implement DP and FL into their ML products. When training an ML model, however, these approaches only apply to some data processing risks — e.g., so that the model can not be used to re-identify data of individuals from the model outputs — and not the full range of risks we discuss in our taxonomy. Owing to these shortcomings, organizations that commonly advocate for end-user privacy rights, like the Electronic Frontier Foundation (EFF), have argued against the use of these approaches when they are used as stand-ins for stronger privacy protections (e.g., as in the case of Google’s attempt to replace third-party browser cookies with “Federated Learning of Cohorts”) (Cyphers, 2021). For example, the “criminality classifier” that takes in photos of people’s faces and claims to predict their likelihood to be a criminal (Wu and Zhang, 2016) could be built with a federated learning architecture. Doing so would not address the physiognomy risk inherent to the idea itself, nor the exclusion and disclosure risks arising from how the data is collected and the inferences shared without consent.
Data Privacy Auditing
Prior work has created data auditing tools, such as the Privacy Meter151515https://github.com/privacytrustlab/ml_privacy_meter (Murakonda and Shokri, 2020), to help practitioners conduct privacy impact assessments on ML models. Doing so allows practitioners to quantify some privacy risks (e.g., membership inference attacks). However, because the Privacy Meter must be applied after the model is trained, it is inherently limited in its ability to mitigate against the risks that arise in the data collection and processing phases of work. In addition, similar to DP and FL, this approach takes a limited view of privacy and only applies to specific data processing risks — e.g., aggregation risks that arise from collective sensitive personal data in the training data.
Ethics Checklists and Toolkits
Prior work in AI ethics has introduced many toolkits to support practitioners in ethical AI development (Wong et al., 2023), some of which also surface privacy risks. For example, Microsoft’s Harms Modeling161616https://learn.microsoft.com/en-us/azure/architecture/guide/responsible-innovation/harms-modeling/ is an activity that includes design exercises and worksheets that help “evaluate potential ways the use of a technology you are building could result in negative outcomes for people and society,” including potential privacy risks. AI ethics checklists such as Deon171717https://github.com/drivendataorg/deon allow practitioners to “add an ethics checklist to [their] data science projects,” which include questions that make practitioners reflect on the collection, storage, and analysis of data containing PII (personally identifiable information). These checklists and toolkits could help practitioners consider a broader range of privacy risks described in our taxonomy (e.g., data collection and dissemination risks). However, these tools approach privacy risks monolithically, and at a high-level (e.g., privacy loss, PII exposure); they provide little guidance to practitioners to consider privacy risks newly created and/or exacerbated by AI (e.g., physiognomy, distortion risks). In other words, the use of such tools relies on practitioners’ individual awareness of AI privacy risks, which prior work has identified as a key barrier to AI privacy work (Lee et al., 2024).
Note that all of these approaches have value and we are not suggesting that they not be used. Rather, we caution against rhetoric that it is possible to create “privacy-preserving” AI/ML technologies using only these approaches.
5.1.1. Future Work: Creating AI-specific privacy guidance
Given that our findings show that AI creates new types of privacy risks and exacerbates existing ones, and that current privacy-preserving AI/ML methods fall short of identifying and addressing many of these risks, there is a need for future work to fill the gap of mitigating privacy risks created and exacerbated by AI. Specifically, our taxonomy opens up a new design space for privacy-preserving AI/ML tools that aim to raise practitioners’ awareness of utility-intrusiveness trade-offs of their AI product ideas (e.g., (Ernala et al., 2021)). For example, prior work in other AI-adjacent fields, such as Robotics, has explored how to correlate desired robot function with a minimally-invasive set of sensors (Eick and Anton, 2020). In the broader context of implementing privacy and security in software products, prior work has found that practitioners still largely see privacy and security in products as an “all or nothing” notion such that privacy comes at the expense of other important objectives (Tahaei et al., 2021; Gutfleisch et al., 2022).
Future work can explore incorporating our AI privacy taxonomy into harm-envisioning techniques, such as Consequence Scanning (doteveryone, 2020), by providing AI privacy risk prompts to capture associated negative consequences holistically. These techniques can help practitioners run lightweight privacy evaluations on AI product ideas, and help them balance the utility and intrusiveness of these products and services across design iterations. With such a tool, we hypothesize that practitioners can better advocate and design for privacy in working contexts that may dissuade this work (Tahaei et al., 2021; Lee et al., 2024).
Our taxonomy can also consolidate promising future research in foregrounding tensions across data pipelines, practices, and stakeholders (i.e., data subjects, data observers, data beneficiaries, and data victims). By mirroring the first step in Rahwan’s Society in the Loop framework (Rahwan, 2018), AI practitioners can make concrete the envisioned value and the stakeholders of their proposed AI concepts. To assist in this process, future work can create artifacts that encourage practitioners to articulate the value proposition of their envisioned product. Based on our taxonomy, then, it may be possible to mine our database for AI privacy incidents about products that are “semantically” similar based on an articulated value proposition. By showing practitioners related AI privacy incidents, they might then be guided to reflect on the utility-intrusiveness trade-off of their envisioned AI product ideas: for whom that value is generated (i.e., data beneficiaries), whose data is processed to unlock that value (i.e., data subjects), who can be impacted by the data pipeline (i.e., data victims), and by which privacy risk (e.g., surveillance).
In practice, however, this type of early-stage discussion around AI utility and privacy risk can be challenging because: (i) practitioners do not necessarily understand the full potential and limitations of AI (Yildirim et al., 2022); (ii) privacy is often treated as compliance with general regulatory mandates rather than a product-specific design choice (Waldman, 2017); and, (iii) practitioners do not have access to AI-specific tools that support their privacy work pertaining to the capabilities and requirements that AI brings to their products (Lee et al., 2024). Accordingly, there is a need for a greater understanding of where such tools and artifacts might be effectively incorporated into practitioners’ workflows.